# Value-Incentivized Preference Optimization: A Unified Approach to Online and Offline RLHF

### v2

@@ 1.1 Our contributions
    In addition, VPO reveals the critical role of reward calibration, where the shift ambiguity of the reward model inherent in the Bradley-Terry model (Bradley and Terry,, 1952) can be exploited to implement additional behavior regularization (Pal et al.,, 2024; Ethayarajh et al.,, 2024). This allows VPO to provide a theoretical foundation for popular conservative offline RL methods (e.g., (Kumar et al.,, 2020)), as well as regularized RLHF methods (e.g., DPOP (Pal et al.,, 2024)).
@@ 1.2 Related work
Since the introduction of the original RLHF framework, there have been many proposed simplifications of the preference alignment procedure and attempts to improve performance, including SLiC (Zhao et al.,, 2023), GSHF (Xiong et al.,, 2023), DPO (Rafailov et al.,, 2023), and its variants, such as Nash-MD (Munos et al.,, 2023), IPO (Azar et al.,, 2024), OAIF (Guo et al.,, 2024), SPO (Swamy et al.,, 2024), GPO (Tang et al.,, 2024), and DPOP (Pal et al.,, 2024). These methods can roughly be grouped into online and offline variants, depending on whether preference data is collected before training (offline) or by using the current policy during training (online).
In offline preference alignment, identity preference optimization (IPO, (Azar et al.,, 2024)) argues that it is problematic to use the Bradley-Terry model in DPO to convert pairwise preferences into pointwise reward values, and proposes an alternative objective function to bypass the use of the Bradley-Terry model. DPO-Positive (DPOP, (Pal et al.,, 2024)) observes a failure mode of DPO that the standard DPO loss can reduce the model’s likelihood on preferred answers, and proposes to add a regularization term to the DPO objective to avoid such a failure mode. On the other hand, online AI feedback (OAIF, (Guo et al.,, 2024)) proposes an online version of DPO, where online preference data from LLM annotators is used to evaluate and update the current LLM policy in an iterative manner. Iterative reasoning preference optimization (Iterative RPO, (Yuanzhe Pang et al.,, 2024)) proposes to add an additional negative log-likelihood term in the DPO loss to improve performances on reasoning tasks. Finally, (Chang et al.,, 2024) proposes to reuse the offline preference data via reset.
@@
The principles of optimism and pessimism are typically implemented via constructing confidence intervals or posterior sampling, which have been demonstrated to be provably efficient in tabular settings (Jin et al.,, 2018; Shi et al.,, 2022). Yet, these approaches have had limited success in conjunction with deep learning architectures (Gawlikowski et al.,, 2021), and many empirical heuristics in turn lack theoretical validation (Kumar et al.,, 2020). VPO draws inspiration from reward-biased exploration (Kumar and Becker,, 1982; Liu et al.,, 2020, 2024; Hung et al.,, 2021; Mete et al.,, 2021) in the standard online RL literature, but significantly broadens its scope to the offline setting and RLHF for the first time.
@@ 3.1 General framework
In view of the sub-optimality of naive MLE for reward estimation (Cesa-Bianchi et al.,, 2017; Rashidinejad et al.,, 2022), and motivated by the effectiveness of reward-biased MLE in online RL (Kumar and Becker,, 1982; Liu et al.,, 2020, 2024), we propose to regularize the reward estimate via
@@ 3.2 Online RLHF: algorithm and theory
The next theorem demonstrates that Algorithm 1 achieves $\widetilde{\mathcal{O}}(\sqrt{T})$ cumulative regret under mild assumptions. The proof is provided in Appendix A. The proof logic follows from that of (Liu et al.,, 2024).
@@ 4 Token-level VPO
Recently, Rafailov et al., (2024) offered an interpretation of DPO using the token-level Markov Decision Process (MDP), aiming at reconciling the gap between the practical fine-tuning of LLMs at the token level and the theoretical formulation of DPO at the sentence level. Fortunately, this interpretation does not require algorithmic modifications. Below we provide a short discussion and highlight that VPO can be similarly interpreted using the token-level MDP.
@@ A.1 Proof of Theorem 1
For ease of presentation, we assume that $\mathcal{R}$ is finite, i.e., $|\mathcal{R}|<\infty$. The general case can be directly obtained using a covering number argument, which we refer to (Liu et al.,, 2024; Jin et al.,, 2022) for interested readers.
@@
The following lemma is adapted from (Liu et al.,, 2024, Proposition 5.3), whose proof is deferred to Appendix A.2.
@@ A.2 Proof of Lemma 2
Lemma 5 ((Zhang,, 2023, Theorem 13.2),(Liu et al.,, 2024, Lemma D.1))

### v3

@@ Abstract
Keywords: Reinforcement learning from human feedback (RLHF), preference optimization, optimism, pessimism
@@ 1.1 Our contributions
    In addition, VPO reveals the critical role of reward calibration, where the shift ambiguity of the reward model inherent in the Bradley-Terry model (Bradley and Terry,, 1952) can be exploited to implement additional behavior regularization (Pal et al.,, 2024; Ethayarajh et al.,, 2024) via centering the reward model with respect to a calibration policy. This allows VPO to provide a theoretical foundation for popular conservative offline RL methods (e.g., (Kumar et al.,, 2020)), as well as regularized RLHF methods (e.g., DPOP (Pal et al.,, 2024)).
@@ 1.2 Related work
Since the introduction of the original RLHF framework, there have been many proposed simplifications of the preference alignment procedure and attempts to improve performance, including but not limited to SLiC (Zhao et al.,, 2023), GSHF (Xiong et al.,, 2023), DPO (Rafailov et al.,, 2023), and its variants, such as Nash-MD (Munos et al.,, 2023), IPO (Azar et al.,, 2024), OAIF (Guo et al.,, 2024), SPO (Swamy et al.,, 2024), SPIN (Chen et al.,, 2024), GPO (Tang et al.,, 2024), and DPOP (Pal et al.,, 2024). These methods can roughly be grouped into online and offline variants, depending on whether preference data is collected before training (offline) or by using the current policy during training (online).
In offline preference alignment, identity preference optimization (IPO, (Azar et al.,, 2024)) argues that it is problematic to use the Bradley-Terry model in DPO to convert pairwise preferences into pointwise reward values, and proposes an alternative objective function to bypass the use of the Bradley-Terry model. DPO-Positive (DPOP, (Pal et al.,, 2024)) observes a failure mode of DPO that the standard DPO loss can reduce the model’s likelihood on preferred answers, and proposes to add a regularization term to the DPO objective to avoid such a failure mode. On the other hand, online AI feedback (OAIF, (Guo et al.,, 2024)) proposes an online version of DPO, where online preference data from LLM annotators is used to evaluate and update the current LLM policy in an iterative manner. Iterative reasoning preference optimization (Iterative RPO, Pang et al., (2024)) proposes to add an additional negative log-likelihood term in the DPO loss to improve performances on reasoning tasks. Finally, Chang et al., (2024) proposes to reuse the offline preference data via reset.
@@
The principles of optimism and pessimism are typically implemented via constructing confidence intervals or posterior sampling, which have been demonstrated to be provably efficient in tabular settings (Jin et al.,, 2018; Shi et al.,, 2022). Yet, these approaches have had limited success in conjunction with deep learning architectures (Gawlikowski et al.,, 2021), and many empirical heuristics in turn lack theoretical validation (Kumar et al.,, 2020). Notwithstanding, alternative regularization schemes are developed for general function approximation settings with theoretical guarantees, such as Bellman-consistent pessimism for offline RL (Xie et al.,, 2021), and reward-biased exploration for online RL (Mete et al.,, 2021; Liu et al., 2024a, ). VPO draws inspiration from reward-biased exploration (Kumar and Becker,, 1982; Liu et al.,, 2020; Liu et al., 2024a, ; Hung et al.,, 2021; Mete et al.,, 2021) in the standard online RL literature, but significantly broadens its scope to the offline setting and RLHF for the first time.
Concurrent work.
Since posting the initial version of this paper on arXiv, we discovered several concurrent work that also appeared online around the same time proposing similar regularization techniques as ours to encourage optimism (resp. pessimism) for online (resp. offline) RLHF (Zhang et al.,, 2024; Xie et al.,, 2024; Liu et al., 2024b, ). Despite slightly different forms, the algorithms studied in Zhang et al., (2024); Xie et al., (2024); Liu et al., 2024b can be interpreted as adopting different choices of the calibration policy in VPO. In the context of online RLHF, Zhang et al., (2024) empirically studied a similar algorithm as the proposed online VPO under the contextual bandit formulation of RLHF; Xie et al., (2024) provided finite-time regret analysis of a similar algorithm for the token-level MDP formulation with general function approximation, which extends to general deterministic contextual MDPs as well. In the context of offline RLHF, Liu et al., 2024b studied a similar algorithm as the proposed offline VPO and provided a sample complexity analysis under the contextual bandit formulation, yet focusing on general function approximation and different assumptions.
@@ 3.1 General framework
In view of the sub-optimality of naive MLE for reward estimation (Cesa-Bianchi et al.,, 2017; Rashidinejad et al.,, 2022), and motivated by the effectiveness of reward-biased MLE in online RL (Kumar and Becker,, 1982; Liu et al.,, 2020; Liu et al., 2024a, ), we propose to regularize the reward estimate via
@@ 3.2 Online RLHF: algorithm and theory
The next theorem demonstrates that Algorithm 1 achieves $\widetilde{\mathcal{O}}(\sqrt{T})$ cumulative regret under mild assumptions. The proof is provided in Appendix A. The proof logic follows from that of (Liu et al., 2024a, ).
@@ 4 Token-level VPO
Recently, Rafailov et al., (2024); Zhong et al., (2024) offered an interpretation of DPO using the token-level Markov Decision Process (MDP), aiming at reconciling the gap between the practical fine-tuning of LLMs at the token level and the theoretical formulation of DPO at the sentence level. Fortunately, this interpretation does not require algorithmic modifications. Below we provide a short discussion and highlight that VPO can be similarly interpreted using the token-level MDP following the setups in Rafailov et al., (2024); Zhong et al., (2024).
@@ A.1 Proof of Theorem 1
For ease of presentation, we assume that $\mathcal{R}$ is finite, i.e., $|\mathcal{R}|<\infty$. The general case can be directly obtained using a covering number argument, which we refer to (Liu et al., 2024a, ; Jin et al.,, 2022) for interested readers.
@@
The following lemma is adapted from (Liu et al., 2024a, , Proposition 5.3), whose proof is deferred to Appendix A.2.
@@ A.2 Proof of Lemma 2
Lemma 5 ((Zhang,, 2023, Theorem 13.2),(Liu et al., 2024a, , Lemma D.1))